A father is 46 years older than his son. In 10 years he'll be twice as old as his son what are their present ages
Answer by stanbon(75887)

A father is 46 years older than his son. In 10 years he'll be twice as old as his son what are their present ages
---
F = S + 46
F+10 = 2(S+10)
-------
Substitute for "F" and solve for "S"::
S + 46 + 10 = 2S + 20
------
S = 36 yrs (son's age now)
---
F = S + 46 = 36+46 = 82 yrs (Father's age now)
----------------
Cheers,
Stan H.
-----------------